### v3

# CRISP-NAM: Competing Risks Interpretable Survival Prediction with Neural Additive Models

Abstract
Competing risks are crucial considerations in survival modelling, particularly in healthcare domains where patients may experience multiple distinct event types. We propose CRISP-NAM (Competing Risks Interpretable Survival Prediction with Neural Additive Models), an interpretable neural additive model for competing risks survival analysis which extends the neural additive architecture to model cause-specific hazards while preserving feature-level interpretability. Each feature contributes independently to risk estimation through dedicated neural networks, allowing for visualization of complex non-linear relationships between covariates and each competing risk. CRISP-NAM demonstrates competitive performance on multiple datasets compared to existing approaches. Code is available at https://github.com/VectorInstitute/crisp-nam
1 Introduction
Survival analysis provides a robust framework for time-to-event prediction across multiple disciplines including medicine, finance, and manufacturing. Traditionally rooted in statistical methods, this analytical approach focuses on using available covariates to predict when specific events will occur. As an example, in healthcare settings, clinicians leverage patient data such as laboratory values, medical history, and comorbidities to forecast hospital readmission timing for specific conditions. Financial institutions similarly could apply these techniques to customer transaction histories and demographic profiles to assess default probability on credit products.
2 Background and Related Work
2.1 Cox Proportional Hazards Model
Historically, the Cox Proportional Hazards (Cox PH model) model has been a popular choice for survival analysis. It is a semi-parametric, linear model that relates covariates to the hazard function, which characterizes the instantaneous risk of an event occurring at time $t$, given survival up to that time. The Cox PH model model assumes that covariates have a multiplicative effect on the hazard and that their effects are constant over time (proportional hazards assumption). Mathematically, the hazard function under this model is expressed as:
$h(t\mid X)=h_{0}(t)\exp(\beta^{\top}X)$ (1)
where $h_{0}(t)$ is the baseline hazard function, $X$ is the vector of covariates, and $\beta$ is the vector of regression coefficients.
Despite its widespread use and interpretability, the Cox PH model has several limitations. For example, nonlinear relationships must be manually specified must be manually specified (e.g., using splines), or alternatively introduced using nonlinear kernels Cai et al. (2011) which can be challenging in high-dimensional settings. Additionally, this model assumes that the effect of each covariate on the hazard is constant over time. Violations of this assumption can lead to biased estimates. Finally, in scenarios with many features or complex interactions, prior feature engineering or dimensionality reduction may be necessary to avoid convergence issues or unstable estimates.
2.2 Deep Learning for Survival Analysis
To address the limitations of traditional statistical models, recent years have seen a shift towards machine learning-based survival models, including neural networks, which can capture nonlinear effects, interactions, and high-dimensional structure in the data.
DeepSurv Katzman et al. (2018) is one of the earliest deep learning models designed for survival analysis. It consists of a deep feedforward network with a single output node with a linear activation which estimates the log-risk function in the Cox PH model. Kvamme et al. (2019) introduced a joint time–covariate network $f(t,x)$, breaking the proportional hazards assumption by modelling the effect of $x$ as varying with time. This is conceptually closer to dynamic hazard models or time-dependent Cox PH models. With the aim of increasing trust and adoption, alignment with medical knowledge and supporting regulatory requirements, researchers have proposed several interpretable survival models in the literature. Kovalev et al. (2020) proposed SURVLIME, which incorporates the Local Interpretable Model-agnostic Explanation (LIME) framework Ribeiro et al. (2016) to approximate the survival model in the local neighbourhood of a test instance in feature space.
Neural Additive Models (NAMs) Agarwal et al. (2021), a neural-network extension of Generalized Additive Models (GAMs) have been used in machine learning based survival models, examples of which are SurvNAM Utkin et al. (2022) and CoxNAM Xu and Guo (2023). While both SurvNAM and CoxNAM employ NAMs Agarwal et al. (2021) to enhance interpretability in survival analysis, they differ fundamentally in purpose and integration. CoxNAM is a fully trainable survival model that embeds NAMs directly within the Cox proportional hazards framework, enabling inherently interpretable, end-to-end learning of nonlinear feature effects from survival data. In contrast, SurvNAM is a post-hoc explanation method that approximates the predictions of a pre-trained black-box survival model such as a Random Survival Forest (RSF) by fitting a GAM-extended Cox PH model using NAMs as surrogate learners.
It is to be noted that none of the survival models discussed thus far are capable of modelling competing risks, which will be covered next.
2.3 Competing Risks in Survival Analysis
While conventional survival models primarily address single outcomes, real-world scenarios frequently involve competing risks that fundamentally alter the probability distribution of the primary event. For instance, if a patient dies, the possibility of experiencing a subsequent heart‑related complication is removed, exemplifying a typical competing‑events situation. Two principal methodological frameworks have emerged for analyzing competing risks:
The Cause-Specific Hazard approach (Prentice et al., 1978) models competing events separately, treating each outcome as a distinct hazard function and censoring subjects who experience competing events from the risk set, without requiring actual independence between event types. For each cause $k$, the cause-specific hazard $\lambda_{k}(t|\mathbf{x})$ represents the instantaneous rate of occurrence of event type $k$ at time $t$ for subjects who have not experienced any event prior to time $t$:
$\small\lambda_{k}(t|\mathbf{x})=\lim_{\Delta t\to 0}\frac{P(t\leq T<t+\Delta t,E=k\mid\mathbf{x})}{\Delta t}$ (2)
where $T$ represents the event time and $E\in\{1,2,\ldots,K\}$ denotes the event type.
In contrast, the Fine-Gray sub-distribution model Fine and Gray (1999) directly accounts for competing events by maintaining subjects who experience competing risks within the risk set. Given the risk set for competing event $k$:
$R_{k}^{sub}(t)=\{i:T_{i}\geq t\text{ or }(T_{i}<t\text{ and }E_{i}\neq k)\}$ (3)
This risk set includes subjects who have either not experienced any event by time $t$ or have experienced a competing event (not event $k$) before time $t$.
The sub-distribution hazard is then defined as:
$\small\lambda_{k}^{sub}(t|\mathbf{x})=\lim_{\Delta t\to 0}\frac{P(t\leq T<t+\Delta t,E=k|i\in R_{k}^{sub}(t),\mathbf{x})}{\Delta t}$ (4)
2.4 Deep Survival Models for Competing Risks
Several deep learning approaches have been proposed to address competing risks in survival analysis. DeepHit Lee et al. (2018) is a joint model for survival analysis with competing risks. It uses a shared representation network followed by cause-specific sub-networks to model the joint distribution of the event time and event type. The model is trained using a combination of the negative log-likelihood and a ranking loss to encourage concordance between predicted risks and observed outcomes. Neural Fine Gray Jeanselme et al. (2023) extends the Fine-Gray sub-distribution model using neural networks to capture non-linear relationships between covariates and sub-distribution hazards. It allows for flexible modelling of competing risks while maintaining the ability to directly estimate cumulative incidence functions.
Despite these advances, a key limitation of existing deep survival approaches for competing risks is their lack of interpretability, especially at the feature level, making it difficult to understand how individual features contribute to risk predictions for different competing events.
To this end, an extension of the Neural Additive Model for competing risks settings is proposed in this paper, resulting in an inherently interpretable survival model. Our approach retains the interpretability and feature-wise transparency of NAMs, while allowing for flexible, non-linear modelling of cause-specific or sub-distribution hazards in competing risks scenarios.
3 Model Architecture
CRISP-NAM extends Neural Additive Models to the competing risks survival analysis setting while preserving feature-level interpretability. The architecture consists of three primary components:
[…]
Figure 1: CRISP-NAM architecture showing the flow from input features through neural networks for two competing risk predictions.
3.1 Neural Additive Model (FeatureNet)
In line with the Neural Additive Model framework, each input feature $x_{i}$ is processed by its own dedicated neural network $f_{i}(\cdot)$, referred to here as a FeatureNet. These feature-specific sub-networks are designed to learn the non-linear contribution of each individual feature to the overall risk score, while preserving interpretability by isolating feature effects.
$\mathbf{h}_{i}=f_{i}(x_{i})\in\mathbb{R}^{d}$ (5)
where $d$ is the dimension of the hidden representation. Each FeatureNet is a fully-connected feedforward neural network with $L$ layers, taking the scalar input $x_{i}$ and producing a hidden representation $\mathbf{h}_{i}\in\mathbb{R}^{d}$. The activations are computed recursively using the hyperbolic tangent function:
$\mathbf{z}^{(l)}_{i}=\tanh(W^{(l)}_{i}\mathbf{z}^{(l-1)}_{i}+b^{(l)}_{i}),\quad\text{for }l=1,\ldots,L,$ (6)
where $\mathbf{z}^{(0)}_{i}=x_{i}$, and $\mathbf{h}_{i}=\mathbf{z}^{(L)}_{i}$ denotes the output of the final layer. Each layer $l$ has weights $W^{(l)}i\in\mathbb{R}^{d_{l}\times d{l-1}}$ and biases $b^{(l)}_{i}\in\mathbb{R}^{d_{l}}$, with $d_{0}=1$ and $d_{L}=d$.
In this implementation, Dropout is used with rate $p_{\text{dropout}}$ after each hidden layer, Feature dropout with rate $p_{\text{feature}}$ during training to increase robustness and an optional batch normalization layer after each linear transformation to stabilize learning especially for deeper FeatureNets.
3.2 Risk-Specific Projections
For each feature $i$ and competing risk $k$, a separate linear projection transforms the feature representation to its contribution to the log-hazard ratio:
$g_{i,k}(\mathbf{h}_{i})=\mathbf{w}_{i,k}^{T}\mathbf{h}_{i}$ (7)
where $\mathbf{w}_{i,k}\in\mathbb{R}^{d}$ is the projection vector for feature $i$ and risk $k$. This risk-specific projection enables the model to capture how the same feature might influence different competing risks differently.
3.3 Additive Risk Aggregation
The cause-specific log-hazard ratio for risk $k$ given input features $\mathbf{x}=[x_{1},x_{2},\ldots,x_{p}]$ is computed as the sum of individual feature contributions:
$\eta_{k}(\mathbf{x})=\sum_{i=1}^{p}g_{i,k}(f_{i}(x_{i}))$ (8)
This preserves the additive nature of the model while allowing for complex non-linear feature effects.
3.4 Cause-Specific Hazards Approach
The cause-specific hazards framework Prentice et al. (1978) is adopted for the CRISP-NAM model. For a subject with covariates $\mathbf{x}$, we parameterize each cause-specific hazard using a Cox-type model:
$\lambda_{k}(t|\mathbf{x})=\lambda_{0k}(t)\exp(\eta_{k}(\mathbf{x}))$ (9)
where $\lambda_{0k}(t)$ is the baseline hazard for the $k$-th event and $\eta_{k}(\mathbf{x})$ is the risk score function for event type $k$. Unlike traditional Cox models with linear risk functions, our approach uses FeatureNets within a neural additive model (NAM), enabling it to model complex non-linear effects of individual features.
3.5 Partial Likelihood Loss Function
To train the model, the standard Cox partial likelihood approach, adapted for competing risks Prentice et al. (1978), is implemented. For a dataset with $n$ subjects, the negative log partial likelihood for event type $k$ is
$\small\mathcal{L}_{k}=-\sum_{i:E_{i}=k}\Bigg{[}\eta_{k}(\mathbf{x}_{i})-\log\left(\sum_{j:T_{j}\geq T_{i}}\exp(\eta_{k}(\mathbf{x}_{j}))\right)\Bigg{]}$ (10)
where $E_{i}$ is the event type for subject $i$ and $T_{i}$ is their event or censoring time. The risk set for each event consists of subjects who have not experienced any event by time $T_{i}$, which is the key characteristic of the cause-specific hazards approach.
The overall loss is the sum of the negative log partial likelihoods across all event types:
$\mathcal{L}=\sum_{k=1}^{K}\mathcal{L}_{k}+\gamma\|\Theta\|^{2}_{2}$ (11)
where $\gamma$ is the $L_{2}$ regularization parameter and $\Theta$ represents all model parameters.
Since many real-world problems involving competing risks suffer from class imbalance, we adopt a risk-frequency-weighted version of the partial likelihood. Specifically, we define:
$\small\mathcal{L}_{k,\omega}=-\omega_{k}\sum_{i:E_{i}=k}\bigg{[}\eta_{k}(\mathbf{x}_{i})-\log\Big{(}\sum_{j:T_{j}\geq T_{i}}\exp\big{(}\eta_{k}(\mathbf{x}_{j})\big{)}\Big{)}\bigg{]}$ (12)
where $\omega_{k}$ is a weight inversely proportional to the frequency of event type $k$. The total loss is given by:
$\mathcal{L}_{\text{weighted}}=\sum_{k}\mathcal{L}_{k,\omega}$ (13)
3.6 Baseline Hazard Estimation
In the proportional hazards formulation of the subdistribution hazard model, the hazard for event type $k$ is expressed as
$\lambda_{k}(t\mid\mathbf{x})=\lambda_{0k}(t)\cdot\exp\big{(}\eta_{k}(\mathbf{x})\big{)}$ (14)
where $\lambda_{0k}(t)$ is the baseline subdistribution hazard function for event type $k$, and $\eta_{k}(\mathbf{x})$ is the covariate-dependent log-risk function produced by the model.
We do not directly parameterize $\lambda_{0k}(t)$; instead, we estimate the corresponding baseline cumulative hazard function $\Lambda_{0k}(t)$ after training using the Breslow estimator Breslow (1975):
$\small\hat{\Lambda}_{0k}(t)=\sum_{i:T_{i}\leq t,E_{i}=k}\frac{1}{\sum_{j:T_{j}\geq T_{i}}\exp\left(\eta_{k}(\mathbf{x}_{j})\right)}$ (15)
This cumulative baseline hazard can then be used to compute the cumulative incidence function (CIF) for each event type $k$ at test time.
Estimating baseline hazards is required in the proposed approach for several reasons. First, while the neural additive component of CRISP-NAM efficiently learns relative risks between subjects (hazard ratios), the baseline hazard estimation enables translation of these relative measures into absolute risk predictions. This is necessary for clinical decision-making, where understanding the actual probability of an event is more informative than just knowing if one patient has higher risk than another. Second, in competing risks settings, accurate baseline hazard estimation is required for proper calculation of cumulative incidence functions, as shown in Eqs. (16) and (17). Third, the baseline hazard captures the underlying temporal pattern of risk independent of covariates, allowing CRISP-NAM to generate time-dependent predictions at clinically relevant horizons (e.g., 1-year, 5-year risks). Finally, proper evaluation metrics like Brier scores and TD-AUC values at specific time points depend on accurate absolute risk estimation.
3.7 Prediction of Absolute Risks
To predict the cumulative incidence function (CIF) for each competing event, we combine the cause-specific hazards approach with the relationship between cause-specific hazards and the CIF. For a subject with covariates $\mathbf{x}$, the CIF for event type $k$ at time $t$ is calculated as
[…]
where $S(t|\mathbf{x})$ is the overall survival function representing the probability of not experiencing any event by time $t$, $u$ is the integration variable representing all potential event times between 0 and $t$, $\lambda_{k}$ is the baseline hazard and $S(u-|\mathbf{x})$ denotes the left limit of the survival function (i.e., the probability of not experiencing any event just before time $u$). Here, $t$ represents the fixed time horizon of interest (e.g., 1-year or 5-year risk), while $u$ captures the continuous nature of when events might occur within that horizon. The overall survival function is defined as
[…]
In practice, a discrete approximation by dividing the time interval into small segments and summing the contributions across these segments using the estimated baseline hazards and the predicted risk scores is used to compute these integrals. Concretely, for discrete time points $\{t_{1},t_{2},\dots,t_{M}\}$, the CIF is approximated by
$\hat{F}_{k}(t|\mathbf{x})\approx\sum_{t_{m}\leq t}\hat{S}(t_{m-1}|\mathbf{x})\cdot\hat{\lambda}_{k}(t_{m}|\mathbf{x})$ (18)
where
$\hat{S}(t_{m-1}|\mathbf{x})=\exp\left(-\sum_{k^{\prime}=1}^{K}\sum_{t_{\ell}<t_{m}}\hat{\lambda}_{k^{\prime}}(t_{\ell}|\mathbf{x})\right)$
and $\hat{\lambda}_{0k}(t_{m})$ is the estimated baseline hazard at time $t_{m}$ for event type $k$.
3.8 Interpretability Mechanisms
Given that CRISP-NAM is based on NAMs, as with all variants of Generalized Additive Models, we can use shape functions plots to visualize the (non-linear) contribution of each feature to the prediction. Specifically, for each feature $i$ and risk $k$, we can extract a shape function that describes how the feature affects the log-hazard ratio.
$s_{i,k}(x_{i})=g_{i,k}(f_{i}(x_{i}))$ (19)
Additionally, the importance of feature $i$ for risk $k$ is quantified by the mean absolute value of its contribution across the dataset.
$\mathcal{I}_{i,k}=\frac{1}{n}\sum_{j=1}^{n}\left|s_{i,k}(x_{ij})\right|$ (20)
This enables ranking features by their impact on each competing risk, providing valuable insights into risk-specific predictor importance.
4 Experiments
In this section, the datasets used and the experimental procedure to evaluate the CRISP-NAM model is described.
4.1 Datasets
In order to evaluate the proposed interpretable model for competing risks survival prediction, we used the following three real-world medical datasets:
Primary Biliary Cholangitis (PBC).
The PBC dataset originates from a randomized controlled trial conducted at the Mayo Clinic between 1974 and 1984, involving 312 patients diagnosed with primary biliary cholangitis. The study aimed to evaluate the efficacy of D-penicillamine in treating the disease. Each patient record includes 25 covariates encompassing demographic, clinical, and laboratory measurements. The primary endpoint was mortality while on the transplant waiting list, with liver transplantation considered a competing risk (Therneau and Grambsch, 2000).
Framingham Heart Study.
Initiated in 1948, the Framingham Heart Study is a longitudinal cohort study designed to investigate cardiovascular disease (CVD) risk factors. For this analysis, data from 4,434 male participants were utilized, each with 18 baseline covariates collected over a 20-year follow-up period. The study focuses on modelling the risk of developing CVD, treating mortality from non-CVD causes as a competing event Kannel and McGee (1979).
SUPPORT2 Dataset.
The SUPPORT2 dataset originates from the Study to Understand Prognoses and Preferences for Outcomes and Risks of Treatments (SUPPORT2), a comprehensive investigation conducted across five U.S. medical centers between 1989 and 1994. This dataset encompasses records of 9,105 critically ill hospitalized adults, each characterized by 42 variables detailing demographic information, physiological measurements, and disease severity indicators. The study was executed in two phases: Phase I (1989–1991) was a prospective observational study aimed at assessing the care and decision-making processes for seriously ill patients, while Phase II (1992–1994) implemented an intervention to enhance end-of-life care. The primary objective was to develop and validate prognostic models estimating 2- and 6-month survival probabilities, thereby facilitating improved clinical decision-making and patient-physician communication regarding treatment preferences and outcomes (Investigators et al., 1995).
Synthetic Dataset.
We use the synthetic dataset introduced by Lee et al. Lee et al. (2018), which models two competing risks with distinct but overlapping covariate effects. Each patient $i$ is assigned a 12-dimensional feature vector $\mathbf{x}^{(i)}\sim\mathcal{N}(0,I_{12})$, partitioned into three 4-dimensional subgroups: $\mathbf{x}_{1}^{(i)},\mathbf{x}_{2}^{(i)},\mathbf{x}_{3}^{(i)}$. The event times are sampled from exponential distributions as:
$T_{1}^{(i)}$ $\sim\exp\left(\gamma_{T}\|\mathbf{x}_{3}^{(i)}\|^{2}+\gamma_{T}\mathbf{1}^{\top}\mathbf{x}_{1}^{(i)}\right),$ (21)
$T_{2}^{(i)}$ $\sim\exp\left(\gamma_{T}\|\mathbf{x}_{3}^{(i)}\|^{2}+\gamma_{T}\mathbf{1}^{\top}\mathbf{x}_{2}^{(i)}\right),$ (22)
where $\gamma_{T}=10$. Covariates $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ influence only their respective event times, while $\mathbf{x}_{3}$ affects both.
The dataset comprises 30,000 rows of unique patient data with 50% random right-censoring by drawing a censoring time $t_{c}^{(i)}\sim\mathcal{U}[0,\min\{T_{1}^{(i)},T_{2}^{(i)}\}]$. The final observed data for each patient is $(\mathbf{x}^{(i)},t^{(i)},k^{(i)})$, where $t^{(i)}$ is the observed time and $k^{(i)}$ is the event indicator ($\emptyset$ if censored).
4.2 Experimental Setup
The CRISP-NAM model is implemented using PyTorch. The experimental procedure is comprised of hyperparameter optimization followed by cross-validation evaluation phases. First, Optuna (Akiba et al., 2019) is employed to systematically search for optimal model configurations, tuning learning rate, $L_{2}$ regularization strength, dropout rates, network architecture (1-3 hidden layers with 32-256 units), and batch normalization settings using validation loss as the objective. The best configuration was then evaluated through stratified 5-fold cross-validation on multiple datasets previously described. Continuous features were normalized using standard scaling ($\mu=0,\sigma=1$). Missing values for categorical variables were imputed using mode imputation, and for continuous variables, mean imputation is applied for all datasets, except for the SUPPORT2 dataset for which median imputation was applied. Training employed the AdamW (Loshchilov and Hutter, 2017) optimizer to minimize the negative log-likelihood loss with a batch size of 256 and early stopping (patience=10) to prevent over-fitting.
Model performance was assessed using complementary metrics (Park et al., 2021) for discrimination and calibration. For discrimination ability, the Time-Dependent Area-Under-the-Curve (TD-AUC) was used to quantify how well the model ranks subjects by risk, with values ranging from 0.5 (no better than chance) to 1.0 (perfect discrimination). Additionally, the Time-Dependent Concordance Index (TD-CI), which considers that the model’s performance can change over time was computed. This is crucial because the risk of an event may evolve as time progresses, and a model’s predictive ability might not be constant. TD-CI ranges from 0.5 to 1.0, with higher values indicating better discriminative ability. The third metric was the Brier score (BS), which measures the accuracy of probabilistic predictions and penalizes both discrimination and calibration errors with lower values of this score indicating better performance. All metrics were evaluated at multiple clinically relevant time horizons corresponding to the 25th, 50th, and 75th percentiles of observed event times for each competing risk.
5 Results and Discussion
Table 1: Comparative 5-fold performance metrics for various models across multiple competing risks datasets.
Dataset Model Risk TD-AUC TD-CI Brier Score
$q_{.25}$ $q_{.50}$ $q_{.75}$ $q_{.25}$ $q_{.50}$ $q_{.75}$ $q_{.25}$ $q_{.50}$ $q_{.75}$
FHS CRISP-NAM 1 0.850±.013 0.834±.009 0.812±.015 0.714±.022 0.712±.019 0.708±.019 0.059±.006 0.111±.004 0.153±.007
2 0.785±.057 0.776±.029 0.769±.037 0.720±.033 0.719±.021 0.714±.021 0.040±.004 0.081±.006 0.130±.007
NFG 1 0.678±.160 0.673±.149 0.666±.133 0.632±.150 0.629±.141 0.628±.134 0.058±.005 0.102±.005 0.134±.007
2 0.617±.126 0.629±.117 0.620±.136 0.612±.156 0.611±.148 0.610±.152 0.041±.003 0.077±.006 0.113±.009
DEEPHIT 1 0.854±.019 0.831±.012 0.807±.009 0.738±.030 0.729±.018 0.724±.017 0.056±.006 0.101±.004 0.134±.004
2 0.796±.053 0.779±.028 0.776±.031 0.737±.030 0.728±.018 0.724±.017 0.038±.003 0.072±.004 0.109±.005
SUP CRISP-NAM 1 0.894±.024 0.839±.041 0.824±.052 0.682±.024 0.644±.017 0.636±.015 0.117±.075 0.216±.106 0.238±.111
2 0.950±.037 0.879±.084 0.834±.107 0.869±.005 0.772±.005 0.734±.006 0.100±.007 0.170±.011 0.215±.018
NFG 1 0.790±.036 0.777±.023 0.797±.020 0.620±.038 0.513±.025 0.438±.025 0.044±.003 0.100±.005 0.115±.005
2 0.902±.004 0.840±.003 0.812±.005 0.847±.008 0.741±.007 0.701±.010 0.092±.004 0.151±.004 0.176±.002
DEEPHIT 1 0.779±.205 0.640±.222 0.665±.188 0.746±.013 0.540±.010 0.497±.011 0.039±.003 0.089±.002 0.101±.003
2 0.955±.025 0.893±.069 0.860±.083 0.863±.008 0.745±.005 0.697±.006 0.087±.005 0.144±.006 0.171±.003
PBC CRISP-NAM 1 0.986±.013 0.960±.029 0.953±.039 0.799±.016 0.783±.013 0.772±.009 0.126±.015 0.155±.016 0.176±.016
2 0.977±.028 0.982±.015 0.985±.013 0.675±.020 0.662±.026 0.637±.019 0.314±.146 0.373±.152 0.388±.153
NFG 1 0.853±.023 0.835±.011 0.824±.027 0.782±.017 0.765±.013 0.756±.016 0.143±.020 0.152±.008 0.170±.013
2 0.491±.055 0.491±.055 0.537±.056 0.227±.022 0.238±.011 0.245±.016 0.092±.054 0.135±.073 0.164±.082
DEEPHIT 1 0.994±.008 0.965±.028 0.959±.034 0.822±.016 0.796±.019 0.776±.013 0.100±.009 0.130±.004 0.152±.011
2 0.978±.036 0.983±.027 0.987±.020 0.728±.024 0.705±.019 0.690±.016 0.037±.006 0.056±.003 0.065±.004
SYN CRISP-NAM 1 0.682±.011 0.675±.007 0.671±.011 0.585±.009 0.591±.005 0.587±.003 0.088±.014 0.188±.034 0.287±.047
2 0.694±.021 0.677±.011 0.670±.009 0.594±.007 0.592±.008 0.586±.007 0.073±.016 0.160±.041 0.246±.059
NFG 1 0.802±.011 0.771±.016 0.717±.016 0.741±.007 0.715±.008 0.670±.006 0.051±.001 0.096±.002 0.161±.003
2 0.815±.013 0.771±.016 0.717±.019 0.739±.024 0.706±.026 0.662±.023 0.049±.002 0.095±.003 0.160±.003
DeepHit 1 0.847±.006 0.833±.010 0.815±.006 0.757±.006 0.734±.007 0.696±.004 0.052±.001 0.097±.002 0.160±.002
2 0.855±.009 0.829±.010 0.810±.012 0.762±.009 0.737±.008 0.695±.007 0.049±.001 0.095±.003 0.159±.001
Notes: Dataset = (FHS: Framingham Heart Study, SUP: SUPPORT2, PBC: Primary Biliary Cirrhosis, SYN: Synthetic); Model= (CRISP-NAM: CRISP-NAM, NFG: Neural Fine Gray, DEEPHIT: DeepHit); Risk = (1: Primary, 2: Competing)
Table 1 displays the 5-fold cross-validated performance of CRISP-NAM against two state-of-the-art neural baselines: DeepHit and Neural Fine Gray (NFG). While DeepHit generally achieves the highest performance, CRISP-NAM demonstrates competitive discrimination with the added benefit of interpretability through its additive structure.
CRISP-NAM achieves discrimination metrics within 2-5% of DeepHit on clinical datasets, demonstrating its competitiveness. On FHS, the TD-AUC gap at $q_{0.25}$ is merely 0.004 (0.850 vs 0.854), with CRISP-NAM actually outperforming DeepHit at $q_{0.50}$ for Risk 1 (0.834 vs 0.831). Similarly, on SUPPORT2 Risk 1, CRISP-NAM exceeds DeepHit’s TD-AUC by substantial margins at $q_{0.25}$ (0.894 vs 0.779) and $q_{0.50}$ (0.839 vs 0.640). For PBC, while DeepHit achieves marginally higher TD-AUC, both models reach near-ceiling performance ($\geq 0.96$), making the practical difference negligible. The synthetic dataset represents the primary challenge, where DeepHit’s flexible architecture captures complex non-linear patterns that CRISP-NAM’s additive structure cannot fully represent. While DeepHit generally achieves lower Brier scores, CRISP-NAM maintains clinically acceptable calibration on most datasets. The calibration gap is smallest on FHS (within 0.01-0.02) and moderate on PBC Risk 1. The larger gaps on SUPPORT2 and PBC Risk 2 reflect the trade-off between interpretability and flexibility in probability estimation. CRISP-NAM exhibits lower variance across folds in several scenarios (e.g., FHS Risk 1: ±0.009 vs ±0.012 for DeepHit at $q_{0.50}$), suggesting more stable predictions.
The performance gaps between CRISP-NAM and DeepHit are often clinically insignificant. A TD-CI difference of 0.02-0.03 rarely changes risk stratification decisions in practice. CRISP-NAM’s competitive performance, combined with its interpretability, positions it as a valuable alternative when model transparency is critical, particularly in high-stakes medical applications requiring explainable predictions.
5.1 Interpretability Analysis
We present shape function plots for the top 10 features per risk as learned by the CRISP-NAM model across 3 datasets: SUPPORT2, Framingham and PBC. Each curve $s_{i,k}(x_{i})$ represents the marginal contribution of feature $x_{i}$ to the log cause-specific hazard for risk $k$. Rug plots beneath each curve illustrate the empirical distribution of feature values, highlighting regions, particularly in the tails, where data are sparse.
Interpretation Note. These shape functions are associational, not causal, and may obscure interactions or time-varying effects. They are estimated under smoothness constraints and can extrapolate in regions with low data density, leading to amplified or flattened effects. As such, the shapes should be used primarily for model validation and debugging. Apparent patterns should be interpreted cautiously, corroborated on external cohorts, and discussed with domain experts before drawing scientific or clinical conclusions. While our primary focus centres on analyzing trends revealed through the shape function plots, we provide limited discussion of the associations between covariates and predicted risks. These discussions serve primarily to demonstrate how our shape plot findings align with or contrast against established medical literature.
5.1.1 Framingham Heart Study Dataset
Figure 2: Shape functions computed with CRISP-NAM model for the top-10 important features: Framingham dataset
Figure 2 presents shape functions from the CRISP-NAM model trained on the Framingham dataset, with separate plots for two competing risks: cardiovascular disease (Risk 1) and non–cardiovascular death (Risk 2). All functions are displayed on the log-hazard scale, where a unit increase of $+0.69$ corresponds to a doubling of the cause-specific hazard.
For Risk 1, AGE, SYSBP, BMI, GLUCOSE, and the binary indicator DIABETES exhibit steadily increasing trends across most of their observed ranges. DIABP and AGE display similar steep upward slopes beyond their respective median values. TOTCHOL shows a mild U-shaped profile with a slight increase at high values. educ has an approximately linear decreasing trend across its range. The binary history flags PREVCHD and PREVAP show negative contributions to the log-hazard.Three data characteristics, rather than a real reversal of risk, explain this result:
  1. 1.
    Sparsity. For binary variables with low prevalence rates (less than 10%) such as PREVAP and PREVCHD, excessive smoothing regularization can distort their true impact on survival outcomes. The shape functions for these rare categorical variables are vulnerable to being inappropriately regressed toward the population mean, causing established cardiovascular risk factors to paradoxically appear protective in the visualized contribution plots. This phenomenon occurs because the limited number of positive cases provides insufficient signal to overcome the model’s smoothing penalties, resulting in misleading shape functions that fail to capture the true elevated risk associated with these clinical conditions.
  2. 2.
    Selection. The original study excluded most people with severe existing heart disease. The retained group is therefore healthier or already under treatment, a “survival‑selection” bias that lowers their short‑term risk estimates D’Agostino and Kannel (1989).
  3. 3.
    Competing‑risk censoring. Deaths due to heart problems are counted under Risk 1. Removing those events from the Risk 1 pool leaves a group that is, by definition, less likely to die from non‑heart causes. This negative effect can then bleed back into the Risk 1 estimate because the true positive effect must be learned from very few events.
For Risk 2, the shape functions for SYSBP and GLUCOSE follow J-shaped profiles, with minimal contribution at lower values and a marked increase beyond the upper quantiles. educ decreases nearly linearly. The binary variable PREVCHD is associated with a negative contribution, while DIABETES presents a discrete step-wise increase. The BMI function shows a shallow U-shape. DIABP increases monotonically throughout its observed range.
Several observed patterns in the shape functions are consistent with known risk factors for cardiovascular and all-cause mortality. Established studies have linked increasing age, elevated systolic blood pressure, high glucose levels, and diabetes with heightened cardiovascular risk (Kannel and McGee, 1979; D’Agostino and Kannel, 1989). The inverse trend for educational attainment aligns with literature on socioeconomic disparities in cardiovascular outcomes (Rosengren et al., 2019). The U-shaped relationship observed for BMI (Risk 2) has been previously noted in older adults and is often described as the “obesity paradox” (Amundson et al., 2010) as well as sparse data in the extreme ranges. Similarly, the positive association between elevated diastolic pressure and non-cardiovascular mortality may reflect constraints due to cohort truncation and sparsity in the extreme blood-pressure ranges.
Figure 3 shows how features contribute positively or negatively to the prediction. While feature importance plots are generally not as informative as shape plots, which can reveal more detailed relationship between covariates, we provide it here for completeness.
Figure 3: Feature Importances computed using CRISP-NAM for the Framingham Dataset.
5.1.2 Primary Biliary Cholangitis (PBC) Dataset
Figure 4: Shape functions computed with CRISP-NAM model for the top-10 important features: PBC dataset
Figure 5: Feature Importances computed using CRISP-NAM for the PBC Dataset.
Figure 4 presents the feature importance and shape functions from the CRISP-NAM model trained on the Primary Biliary Cholangitis (PBC) dataset. The plots show distinct patterns for Risk 1 (death on the waiting list) and Risk 2 (transplantation). Additionally, Figure 5 shows top 5 positive and top 5 negative important features that contributed to the model’s prediction.
For Risk 1, age exhibits a monotonic increasing contribution to the predicted hazard. SGOT and alkaline display sharp increases followed by plateaus at higher values. The function for serBilir rises steeply at low values and saturates thereafter. serChol peaks at moderate levels and decreases subsequently. Binary indicators such as hepatomegaly_Yes and edema_edema despite diuretics show positive contributions. albumin demonstrates an inverted S-shaped curve: initially increasing, crossing zero near the median, and declining at higher values. platelet decreases steadily, with a minor upturn in the extreme right tail where data are sparse.
For Risk 2, the shape for age decreases monotonically. SGOT rises rapidly and then flattens. histologic shows an approximately linear increasing trend. The shape for serBilir forms an inverted-U. albumin rises beyond the cohort median and levels off at higher values. ascites_Yes and hepatomegaly_Yes are associated with negative contributions. alkaline increases sharply before flattening. spiders_Yes exhibits a slight negative effect. Finally, serChol increases with higher values, especially in the upper tail, although data in that region are sparse.
The rug plots accompanying each shape function reflect the distribution of the feature values and indicate regions with limited data support.These empirical patterns align with several well-established clinical insights in the context of Primary Biliary Cholangitis (PBC). For instance, older age, elevated liver enzymes such as SGOT and alkaline phosphatase, and increased serum bilirubin are recognized markers of disease severity and poorer prognosis (Dickson et al., 1989; Murillo Perez et al., 2023). The decreasing transplant hazard with age may reflect clinical prioritization criteria that favour younger candidates for organ allocation. The non-linear shape for albumin aligns with its known role as a proxy for liver synthetic function, where low levels indicate hepatic decompensation. Histologic stage progression, from fibrosis to cirrhosis, is a standard determinant in transplant eligibility, consistent with the monotonic rise observed in its shape function. Additionally, ascites and hepatomegaly are classical signs of advanced liver disease, often associated with higher mortality and reduced transplant suitability. The distinct inverted-U pattern for bilirubin in the transplant hazard may reflect clinical decision-making thresholds, while the rising contribution of cholesterol could be interpreted in light of its association with bile duct dysfunction in PBC.
5.1.3 SUPPORT2 Dataset
Figure 6: Shape functions computed with CRISP-NAM model for the top-10 important features: SUPPORT2 dataset
Figure 6 presents feature importance and shape functions from the CRISP-NAM model trained on the SUPPORT2 dataset, which distinguishes cancer-specific mortality (Risk 1) from death due to other causes (Risk 2). All shape functions are plotted on the log-hazard scale.
For Risk 1, the variable sps shows a near-linear increase, indicating a progressive rise in the predicted hazard across its range. The shape function for age follows an inverted-U profile, with risk increasing up to approximately 65 years and declining thereafter. Binary indicators dementia_1 and dnr_no_dnr exhibit negative contributions to the log-hazard. The shape for avtisst is flat at lower durations of mechanical ventilation but shows a steep increase beyond approximately 3 days. Negative contributions for dnr_no_dnr and dementia_1 may reflect treatment-limiting decisions or documentation timing effects, as suggested in earlier SUPPORT studies (Hakim et al., 1996).
For Risk 2, the severity scores aps and sps both show sharp increases followed by plateaus. The avtisst variable again displays a marked rise in hazard for durations exceeding 3 days. Cost-related variables charges and totcst show decreasing trends. The binary indicator dnr_no_dnr also shows a downward slope across its two levels. Rising sps and aps scores are consistent with the role of physiological instability and organ dysfunction in predicting mortality (Knaus et al., 1995). The inverted-U for age in cancer mortality likely reflects competing risks from other causes in older individuals (Hayes-Larson et al., 2020). The association between prolonged ventilation (avtisst $>3$ days) and increased hazard aligns with the known severity of illness in patients requiring extended respiratory support. Cost variables likely act as proxies for length of stay or illness trajectory rather than direct predictors.
Figure 7: Feature Importances computed using CRISP-NAM for the SUPPORT2 Dataset.
6 Conclusion
We introduce CRISP-NAM, a deep survival model that simultaneously addresses competing risks in survival analysis while remaining inherently interpretable. Our model demonstrates competitive discriminative performance on real-world clinical data and uniquely reveals covariate effects through intuitive shape function plots. CRISP-NAM is particularly valuable in high-stakes healthcare ML applications requiring mechanistic understanding such as investigating associational relationships, assessing treatment efficacy, or designing targeted interventions, especially when competing events represent distinct clinical processes (Austin et al., 2016).

### v5

@@
 CRISP-NAM demonstrates competitive performance on multiple datasets compared to existing approaches.
-Code is available at https://github.com/VectorInstitute/crisp-nam
 1 Introduction
@@
 Traditionally rooted in statistical methods, this analytical approach focuses on using available covariates to predict when specific events will occur.
-As an example, in healthcare settings, clinicians leverage patient data such as laboratory values, medical history, and comorbidities to forecast hospital readmission timing for specific conditions.
-Financial institutions similarly could apply these techniques to customer transaction histories and demographic profiles to assess default probability on credit products.
+As an example, in healthcare settings, clinicians leverage patient data such as laboratory values, medical history, and comorbidities to forecast hospital readmission timing for specific medical conditions.
+Many real-world survival scenarios often involve competing risks where multiple mutually exclusive events can occur.
+While recent deep learning advances have improved predictive performance in competing risks settings, these models operate as “black boxes”.
+This opacity poses significant challenges in high-stakes domains like healthcare, where understanding feature contribution to a model’s predictions is crucial for clinical decision-making, regulatory compliance, and building trust with practitioners.
+Moreover, multiple jurisdictions have established interpretability requirements across general and sector-specific regulations.
+For example, Canada’s Artificial Intelligence and Data Act (AIDA) (Government of Canada, 2024) emphasizes risk-based governance with interpretability assessments during development phases.
+Sector-specific regulations include Food and Drug Administration (FDA) guidance (U.S.
+Food and Drug Administration, 2024) requiring clear documentation of AI/ML medical device decision-making and Article 13 of the EU AI Act (art, 2024) which requires high risk AI systems to be designed with sufficient transparency and technical capabilities to explain their outputs, accompanied by clear instructions that enable users to properly interpret and appropriately use the system.
+In this work, we bridge a critical gap between interpretability and competing risks modelling in deep survival analysis for healthcare applications by introducing CRISP-NAM (Competing Risks Interpretable Survival Prediction with Neural Additive Models).
+The main contributions of our work are as follows:
+  • We extend Neural Additive Models (NAMs) to competing risks settings, whereas existing NAM-based survival models only handle single-event outcomes.
+  • We introduce separate projection functions for each risk-feature pair, allowing features to have differential effects across competing events while maintaining interpretability.
+  • The proposed model jointly learns all cause-specific hazards in a single model, rather than treating competing events as censoring or requiring separate models.
+  • Our model can be used to generate shape functions and feature importance rankings for each competing risk and this would allow practitioners to understand how different covariates influence each outcome.
+  • We incorporate risk-frequency weightings to address class imbalances in competing events which is a common challenge that appears in real-world medical datasets.
+The next section reviews the relevant background and related work that motivates our proposed approach.
 2 Background and Related Work
 2.1 Cox Proportional Hazards Model
-Historically, the Cox Proportional Hazards (Cox PH model) model has been a popular choice for survival analysis.
+Historically, the Cox Proportional Hazards (Cox PH) model (Cox, 1972) has been a popular choice for survival analysis.
 It is a semi-parametric, linear model that relates covariates to the hazard function, which characterizes the instantaneous risk of an event occurring at time $t$, given survival up to that time.
-The Cox PH model model assumes that covariates have a multiplicative effect on the hazard and that their effects are constant over time (proportional hazards assumption).
+The Cox PH model assumes that covariates have a multiplicative effect on the hazard and that their effects are constant over time (proportional hazards assumption).
 Mathematically, the hazard function under this model is expressed as:
@@
 Despite its widespread use and interpretability, the Cox PH model has several limitations.
-For example, nonlinear relationships must be manually specified must be manually specified (e.g., using splines), or alternatively introduced using nonlinear kernels Cai et al.
+For example, nonlinear relationships must be manually specified (e.g., using splines), or alternatively introduced using nonlinear kernels Cai et al.
 (2011) which can be challenging in high-dimensional settings.
@@
 CoxNAM is a fully trainable survival model that embeds NAMs directly within the Cox proportional hazards framework, enabling inherently interpretable, end-to-end learning of nonlinear feature effects from survival data.
-In contrast, SurvNAM is a post-hoc explanation method that approximates the predictions of a pre-trained black-box survival model such as a Random Survival Forest (RSF) by fitting a GAM-extended Cox PH model using NAMs as surrogate learners.
-It is to be noted that none of the survival models discussed thus far are capable of modelling competing risks, which will be covered next.
+In contrast, SurvNAM is a post-hoc explanation method that approximates the predictions of a pre-trained black-box survival model such as a Random Survival Forest (Ishwaran et al., 2008) by fitting a GAM-extended Cox PH model using NAMs as surrogate learners.
+Notably, none of the survival models discussed thus far are capable of modelling competing risks, which will be covered next.
 2.3 Competing Risks in Survival Analysis
-While conventional survival models primarily address single outcomes, real-world scenarios frequently involve competing risks that fundamentally alter the probability distribution of the primary event.
-For instance, if a patient dies, the possibility of experiencing a subsequent heart‑related complication is removed, exemplifying a typical competing‑events situation.
-Two principal methodological frameworks have emerged for analyzing competing risks:
+While conventional survival models primarily address single outcomes, many real-world scenarios involve competing risks that fundamentally alter the probability distribution of the primary event.
+For instance, if a patient dies, the possibility of experiencing a subsequent heart‑related complication is removed, illustrating a typical competing‑events situation.
+Two methodological frameworks have emerged for analyzing competing risks:
 The Cause-Specific Hazard approach (Prentice et al., 1978) models competing events separately, treating each outcome as a distinct hazard function and censoring subjects who experience competing events from the risk set, without requiring actual independence between event types.
@@
 $\small\lambda_{k}(t|\mathbf{x})=\lim_{\Delta t\to 0}\frac{P(t\leq T<t+\Delta t,E=k\mid\mathbf{x})}{\Delta t}$ (2)
-where $T$ represents the event time and $E\in\{1,2,\ldots,K\}$ denotes the event type.
+where $\mathbf{x}$ represents the covariates, $T$ represents the event time and $E\in\{1,2,\ldots,K\}$ denotes the event type.
 In contrast, the Fine-Gray sub-distribution model Fine and Gray (1999) directly accounts for competing events by maintaining subjects who experience competing risks within the risk set.
 Given the risk set for competing event $k$:
-$R_{k}^{sub}(t)=\{i:T_{i}\geq t\text{ or }(T_{i}<t\text{ and }E_{i}\neq k)\}$ (3)
-This risk set includes subjects who have either not experienced any event by time $t$ or have experienced a competing event (not event $k$) before time $t$.
+$R_{k}^{sub}(t)=\{j:T_{j}\geq t\text{ or }(T_{j}<t\text{ and }E_{j}\neq k)\}$ (3)
+This risk set includes subjects, $j$, who have either not experienced any event by time $t$ or have experienced a competing event (not event $k$) before time $t$.
 The sub-distribution hazard is then defined as:
-$\small\lambda_{k}^{sub}(t|\mathbf{x})=\lim_{\Delta t\to 0}\frac{P(t\leq T<t+\Delta t,E=k|i\in R_{k}^{sub}(t),\mathbf{x})}{\Delta t}$ (4)
+$\small\lambda_{k}^{sub}(t|\mathbf{x})=\lim_{\Delta t\to 0}\frac{P(t\leq T<t+\Delta t,E=k|j\in R_{k}^{sub}(t),\mathbf{x})}{\Delta t}$ (4)
 2.4 Deep Survival Models for Competing Risks
-Several deep learning approaches have been proposed to address competing risks in survival analysis.
+Deep Survival Models leverage deep learning techniques to address competing risks in survival analysis, offering the ability to model complex non-linear patterns in risk prediction.
 DeepHit Lee et al.
@@
 Despite these advances, a key limitation of existing deep survival approaches for competing risks is their lack of interpretability, especially at the feature level, making it difficult to understand how individual features contribute to risk predictions for different competing events.
+In Fig.
+1, we depict the current gap in the literature of deep survival models.
+Specifically, we are interested in these criteria: Interpretability, Non-Linear Modelling and Competing Risks Capability.
+Models such as SurvNAM and CoxNAM are interpretable, however, they have not been reported to be used in competing risks settings.
+In contrast, the Neural Fine-Gray and DeepHit architectures can model competing risks, but are “black-box” models and can only be explained using post-hoc explainability methods such as SHAP and Partial Dependence Plots.
+Post hoc methods are known to be imprecise and can generate misleading explanations Huang and Marques-Silva (2024); Laugel et al.
+(2019).
+The original formulations of Fine-Gray and Cox PH models are incapable of modelling non-linear relationships.
+Our proposed CRISP-NAM model addresses these gaps in survival models fulfilling all 3 criteria while providing competitive performance.
+Figure 1: Venn diagram situating CRISP-NAM against existing deep survival methods
 To this end, an extension of the Neural Additive Model for competing risks settings is proposed in this paper, resulting in an inherently interpretable survival model.
-Our approach retains the interpretability and feature-wise transparency of NAMs, while allowing for flexible, non-linear modelling of cause-specific or sub-distribution hazards in competing risks scenarios.
+Our approach retains the interpretability and feature-wise transparency of NAMs and allowing for flexible, non-linear modelling of cause-specific or sub-distribution hazards in competing risks scenarios.
 3 Model Architecture
@@
 3.2 Risk-Specific Projections
-For each feature $i$ and competing risk $k$, a separate linear projection transforms the feature representation to its contribution to the log-hazard ratio:
-$g_{i,k}(\mathbf{h}_{i})=\mathbf{w}_{i,k}^{T}\mathbf{h}_{i}$ (7)
-where $\mathbf{w}_{i,k}\in\mathbb{R}^{d}$ is the projection vector for feature $i$ and risk $k$.
-This risk-specific projection enables the model to capture how the same feature might influence different competing risks differently.
+For each feature $i$ and competing risk $k$, a separate linear projection transforms the feature representation to its contribution to the log-hazard ratio.
+To address scale ambiguities across different competing risks and ensure fair comparison of feature contributions, we constrain the projection vectors to have unit L2 norm.
+The risk-specific projection is defined as:
+$g_{i,k}(\mathbf{h}_{i})=\tilde{\mathbf{w}}_{i,k}^{T}\mathbf{h}_{i}$ (7)
+where $\tilde{\mathbf{w}}_{i,k}$ is the L2-normalized projection vector:
+$\tilde{\mathbf{w}}_{i,k}=\frac{\mathbf{w}_{i,k}}{\|\mathbf{w}_{i,k}\|_{2}+\epsilon}$ (8)
+with $\mathbf{w}_{i,k}\in\mathbb{R}^{d}$ being the learnable weight vector for projection $i\in\{1,2,\ldots,p\}$ and risk $k\in\{1,2,\ldots,K\}$, and $\mathbf{h}_{i}\in\mathbb{R}^{d}$ being the feature representation.
+This normalization constraint ensures that $\|\tilde{\mathbf{w}}_{i,k}\|_{2}=1$ for all feature-risk pairs, which constraints all projection vectors to operate on the same scale and enabling direct comparison of feature importance across different competing risks.
 3.3 Additive Risk Aggregation
@@
 (1978), is implemented.
-For a dataset with $n$ subjects, the negative log partial likelihood for event type $k$ is
-$\small\mathcal{L}_{k}=-\sum_{i:E_{i}=k}\Bigg{[}\eta_{k}(\mathbf{x}_{i})-\log\left(\sum_{j:T_{j}\geq T_{i}}\exp(\eta_{k}(\mathbf{x}_{j}))\right)\Bigg{]}$ (10)
-where $E_{i}$ is the event type for subject $i$ and $T_{i}$ is their event or censoring time.
-The risk set for each event consists of subjects who have not experienced any event by time $T_{i}$, which is the key characteristic of the cause-specific hazards approach.
+For a dataset with $N$ subjects, the negative log partial likelihood for event type $k\in\{1,\dots,K\}$ is
+$\small\mathcal{L}_{k}=-\sum_{\begin{subarray}{c}n=1\\ E_{n}=k\end{subarray}}^{N}\Bigg[\eta_{k}(\mathbf{x}_{n})-\log\left(\sum_{\begin{subarray}{c}j=1\\ T_{j}\geq T_{n}\end{subarray}}^{N}\exp(\eta_{k}(\mathbf{x}_{j}))\right)\Bigg]$ (11)
+where $E_{n}$ is the event type for subject $n$ and $T_{n}$ is their event or censoring time.
+The risk set at time $T_{n}$ consists of all subjects $j$ who have not yet experienced any event ($T_{j}\geq T_{n}$), and $\mathbf{x}_{j}$ denotes the feature vector for each subject $j$ in this risk set.
 The overall loss is the sum of the negative log partial likelihoods across all event types:
@@
 where $\gamma$ is the $L_{2}$ regularization parameter and $\Theta$ represents all model parameters.
-Since many real-world problems involving competing risks suffer from class imbalance, we adopt a risk-frequency-weighted version of the partial likelihood. Specifically, we define:
-$\small\mathcal{L}_{k,\omega}=-\omega_{k}\sum_{i:E_{i}=k}\bigg{[}\eta_{k}(\mathbf{x}_{i})-\log\Big{(}\sum_{j:T_{j}\geq T_{i}}\exp\big{(}\eta_{k}(\mathbf{x}_{j})\big{)}\Big{)}\bigg{]}$ (12)
+Since many real-world problems involving competing risks suffer from class imbalance, we adopt a risk-frequency-weighted version of the partial likelihood.
+Specifically, we define:
+$\small\mathcal{L}_{k,\omega}=-\omega_{k}\sum_{\begin{subarray}{c}n=1\\ E_{n}=k\end{subarray}}^{N}\bigg[\eta_{k}(\mathbf{x}_{n})-\log\Big(\sum_{\begin{subarray}{c}j=1\\ T_{j}\geq T_{n}\end{subarray}}^{N}\exp\big(\eta_{k}(\mathbf{x}_{j})\big)\Big)\bigg]$ (13)
 where $\omega_{k}$ is a weight inversely proportional to the frequency of event type $k$. The total loss is given by:
-$\mathcal{L}_{\text{weighted}}=\sum_{k}\mathcal{L}_{k,\omega}$ (13)
+$\mathcal{L}_{\text{weighted}}=\sum_{k=1}^{K}\mathcal{L}_{k,\omega}$ (14)
 3.6 Baseline Hazard Estimation
-In the proportional hazards formulation of the subdistribution hazard model, the hazard for event type $k$ is expressed as
-$\lambda_{k}(t\mid\mathbf{x})=\lambda_{0k}(t)\cdot\exp\big{(}\eta_{k}(\mathbf{x})\big{)}$ (14)
-where $\lambda_{0k}(t)$ is the baseline subdistribution hazard function for event type $k$, and $\eta_{k}(\mathbf{x})$ is the covariate-dependent log-risk function produced by the model.
-We do not directly parameterize $\lambda_{0k}(t)$; instead, we estimate the corresponding baseline cumulative hazard function $\Lambda_{0k}(t)$ after training using the Breslow estimator Breslow (1975):
-$\small\hat{\Lambda}_{0k}(t)=\sum_{i:T_{i}\leq t,E_{i}=k}\frac{1}{\sum_{j:T_{j}\geq T_{i}}\exp\left(\eta_{k}(\mathbf{x}_{j})\right)}$ (15)
-This cumulative baseline hazard can then be used to compute the cumulative incidence function (CIF) for each event type $k$ at test time.
-Estimating baseline hazards is required in the proposed approach for several reasons.
-First, while the neural additive component of CRISP-NAM efficiently learns relative risks between subjects (hazard ratios), the baseline hazard estimation enables translation of these relative measures into absolute risk predictions.
-This is necessary for clinical decision-making, where understanding the actual probability of an event is more informative than just knowing if one patient has higher risk than another.
-Second, in competing risks settings, accurate baseline hazard estimation is required for proper calculation of cumulative incidence functions, as shown in Eqs.
+In the cause-specific proportional hazards formulation, the hazard for event type $k\in\{1,\dots,K\}$ at time $t\geq 0$ for a subject with covariates $\mathbf{x}\in\mathbb{R}^{p}$ is expressed as
+$\lambda_{k}(t\mid\mathbf{x})=\lambda_{0k}(t)\cdot\exp\!\big(\eta_{k}(\mathbf{x})\big),$ (15)
+where
+  • $\lambda_{0k}(t)$ is the baseline cause-specific hazard function for event type $k$, independent of covariates,
+  • $\eta_{k}(\mathbf{x})$ is the covariate-dependent log-risk function produced by the model.
+We do not directly parameterize $\lambda_{0k}(t)$.
+Instead, we estimate the corresponding baseline cumulative hazard function $\lambda_{0k}(t)=\int_{0}^{t}\lambda_{0k}(u)\,du$ after training using the Breslow estimator Breslow (1975):
+$\hat{\lambda}_{0k}(t_{m})=\sum_{\begin{subarray}{c}n:T_{n}\leq t\\ E_{n}=k\end{subarray}}\frac{1}{\sum_{j:T_{j}\geq T_{i}}\exp\!\left(\eta_{k}(\mathbf{x}_{j})\right)},$ (16)
+where $T_{n}$ denotes the observed time for subject $n$, $E_{n}\in\{0,1,\dots,K\}$ is the event indicator ($E_{n}=0$ if censored), and the denominator represents the sum of relative risks for all subjects still at risk at time $T_{n}$.
+This estimated baseline cumulative hazard $\hat{\lambda}_{0k}(t_{m})$ can then be used to compute the cumulative incidence function (CIF) for each event type $k$ at test time.
+Estimating baseline hazards is necessary for several reasons.
+First, while the neural additive component of CRISP-NAM efficiently learns relative risks between subjects (hazard ratios), baseline hazard estimation enables translation of these relative measures into absolute risk predictions.
+This is required for clinical decision-making, where probabilities of events are needed.
+Second, in competing risks settings, accurate baseline hazard estimation is required for proper calculation of CIFs, as shown in Eqs.
 (17) and (18).
 Third, the baseline hazard captures the underlying temporal pattern of risk independent of covariates, allowing CRISP-NAM to generate time-dependent predictions at clinically relevant horizons (e.g., 1-year, 5-year risks).
-Finally, proper evaluation metrics like Brier scores and TD-AUC values at specific time points depend on accurate absolute risk estimation.
+Finally, proper evaluation metrics such as Brier scores and time-dependent AUCs at specific time points depend on accurate absolute risk estimation.
 3.7 Prediction of Absolute Risks
-To predict the cumulative incidence function (CIF) for each competing event, we combine the cause-specific hazards approach with the relationship between cause-specific hazards and the CIF.
-For a subject with covariates $\mathbf{x}$, the CIF for event type $k$ at time $t$ is calculated as
-[…]
-where $S(t|\mathbf{x})$ is the overall survival function representing the probability of not experiencing any event by time $t$, $u$ is the integration variable representing all potential event times between 0 and $t$, $\lambda_{k}$ is the baseline hazard and $S(u-|\mathbf{x})$ denotes the left limit of the survival function (i.e., the probability of not experiencing any event just before time $u$).
-Here, $t$ represents the fixed time horizon of interest (e.g., 1-year or 5-year risk), while $u$ captures the continuous nature of when events might occur within that horizon.
-The overall survival function is defined as
-[…]
-In practice, a discrete approximation by dividing the time interval into small segments and summing the contributions across these segments using the estimated baseline hazards and the predicted risk scores is used to compute these integrals.
-Concretely, for discrete time points $\{t_{1},t_{2},\dots,t_{M}\}$, the CIF is approximated by
-$\hat{F}_{k}(t|\mathbf{x})\approx\sum_{t_{m}\leq t}\hat{S}(t_{m-1}|\mathbf{x})\cdot\hat{\lambda}_{k}(t_{m}|\mathbf{x})$ (18)
+To predict the cumulative incidence function (CIF) (Gray, 1988) for each competing event, we use the relationship between cause-specific hazards and the CIF.
+For a subject with covariates $\mathbf{x}$, the CIF for event type $k$ at time $t\geq 0$ is defined as
+$F_{k}(t|\mathbf{x})=\int_{0}^{t}S(u|\mathbf{x})\,\lambda_{k}(u|\mathbf{x})\,du,$ (17)
 where
-$\hat{S}(t_{m-1}|\mathbf{x})=\exp\left(-\sum_{k^{\prime}=1}^{K}\sum_{t_{\ell}<t_{m}}\hat{\lambda}_{k^{\prime}}(t_{\ell}|\mathbf{x})\right)$
-and $\hat{\lambda}_{0k}(t_{m})$ is the estimated baseline hazard at time $t_{m}$ for event type $k$.
+  • $S(t|\mathbf{x})$ is the overall survival function, i.e., the probability of not experiencing any event up to time $t$,
+  • $u$ is the integration variable representing time between 0 and $t$,
+  • $\lambda_{k}(u|\mathbf{x})=\lambda_{0k}(u)\exp(\eta_{k}(\mathbf{x}))$ is the cause-specific hazard for cause $k$.
+The survival function is given by
+$S(t|\mathbf{x})=\exp\!\left(-\sum_{k=1}^{K}\int_{0}^{t}\lambda_{k}(u|\mathbf{x})\,du\right).$ (18)
+In practice, a discrete approximation is used to compute these integrals.
+Let $\{t_{1},t_{2},\dots,t_{M}\}$ denote a set of ordered discrete time points with $t_{m}\in[0,t]$.
+Then the CIF can be approximated as
+$\hat{F}_{k}(t|\mathbf{x})\approx\sum_{t_{m}\leq t}\hat{S}(t_{m-1}|\mathbf{x})\cdot\hat{\lambda}_{k}(t_{m}|\mathbf{x}),$ (19)
+where
+$\hat{S}(t_{m-1}|\mathbf{x})$ $=\exp\!\left(-\sum_{k^{\prime}=1}^{K}\sum_{t_{\ell}<t_{m}}\hat{\lambda}_{k^{\prime}}(t_{\ell}|\mathbf{x})\right),$
+$\hat{\lambda}_{k}(t_{m}|\mathbf{x})$ $=\hat{\lambda}_{0k}(t_{m})\exp\!\big(\eta_{k}(\mathbf{x})\big),$
+and $\hat{\lambda}_{0k}(t_{m})$ denotes the estimated baseline cause-specific hazard at time $t_{m}$ for event type $k$.
 3.8 Interpretability Mechanisms
-Given that CRISP-NAM is based on NAMs, as with all variants of Generalized Additive Models, we can use shape functions plots to visualize the (non-linear) contribution of each feature to the prediction.
+Given that CRISP-NAM is based on NAMs, as with all variants of Generalized Additive Models, CRISP-NAM can generate shape functions plots to visualize the (non-linear) contribution of each feature to the prediction.
 Specifically, for each feature $i$ and risk $k$, we can extract a shape function that describes how the feature affects the log-hazard ratio.
 $s_{i,k}(x_{i})=g_{i,k}(f_{i}(x_{i}))$ (20)
-Additionally, the importance of feature $i$ for risk $k$ is quantified by the mean absolute value of its contribution across the dataset.
+The importance of feature $i$ for risk $k$ is quantified by the mean absolute value of its contribution across the dataset.
 $\mathcal{I}_{i,k}=\frac{1}{N}\sum_{j=1}^{N}\left|s_{i,k}(x_{ij})\right|$ (21)
 This enables ranking features by their impact on each competing risk, providing valuable insights into risk-specific predictor importance.
+Notably, in this current implementation, CRISP-NAM does not capture features interactions.
+This is by design to prioritize interpretability through independent feature level shape functions, ensuring that feature contributions can be visualized and understood in isolation.
+Adding separate FeatureNets to model feature interactions adds to the model complexity and affects its interpretability as visualization beyond pairwise features interactions is challenging.
+With $p$ features, there are $\mathcal{O}(p^{2})$ possible pairwise interactions, and deciding which interactions to model would also require domain knowledge.
 4 Experiments
-In this section, the datasets used and the experimental procedure to evaluate the CRISP-NAM model is described.
+In this section, the datasets used and the experimental procedure to evaluate the CRISP-NAM model are described.
 4.1 Datasets
-In order to evaluate the proposed interpretable model for competing risks survival prediction, we used the following three real-world medical datasets:
+In order to evaluate the proposed interpretable model for competing risks survival prediction, we used the following three real-world medical datasets and a synthetic dataset.
+Table 1 provides a summary and breakdown of the primary and competing risks for each the datasets used in this work.
 Primary Biliary Cholangitis (PBC).
@@
 Each patient record includes 25 covariates encompassing demographic, clinical, and laboratory measurements.
-The primary endpoint was mortality while on the transplant waiting list, with liver transplantation considered a competing risk (Therneau and Grambsch, 2000).
+The primary endpoint was mortality while on the transplant waiting list, with liver transplantation considered a competing risk (Dickson et al., 1989).
 Framingham Heart Study.
@@
 This dataset encompasses records of 9,105 critically ill hospitalized adults, each characterized by 42 variables detailing demographic information, physiological measurements, and disease severity indicators.
-The study was executed in two phases: Phase I (1989–1991) was a prospective observational study aimed at assessing the care and decision-making processes for seriously ill patients, while Phase II (1992–1994) implemented an intervention to enhance end-of-life care.
-The primary objective was to develop and validate prognostic models estimating 2- and 6-month survival probabilities, thereby facilitating improved clinical decision-making and patient-physician communication regarding treatment preferences and outcomes (Investigators et al., 1995).
+The study was executed in two phases: Phase I (1989–1991) was a prospective observational study aimed at assessing the care and decision-making processes for seriously ill patients and Phase II (1992–1994) implemented an intervention to enhance end-of-life care.
+The primary objective was to develop and validate prognostic models estimating 2- and 6-month survival probabilities, thereby facilitating improved clinical decision-making and patient-physician communication regarding treatment preferences and outcomes (sup, 1995).
 Synthetic Dataset.
@@
 where $\gamma_{T}=10$. Covariates $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ influence only their respective event times, while $\mathbf{x}_{3}$ affects both.
-The dataset comprises 30,000 rows of unique patient data with 50% random right-censoring by drawing a censoring time $t_{c}^{(i)}\sim\mathcal{U}[0,\min\{T_{1}^{(i)},T_{2}^{(i)}\}]$.
+The dataset consists of 30,000 rows of unique patient data with 50% random right-censoring by drawing a censoring time $t_{c}^{(i)}\sim\mathcal{U}[0,\min\{T_{1}^{(i)},T_{2}^{(i)}\}]$.
 The final observed data for each patient is $(\mathbf{x}^{(i)},t^{(i)},k^{(i)})$, where $t^{(i)}$ is the observed time and $k^{(i)}$ is the event indicator ($\emptyset$ if censored).
+Table 1: Dataset characteristics and competing risk statistics
+Dataset Observations Features Primary Competing Risk Censored
+PBC 312 25 Death (44.87%) Transplant (9.29%) 45.83%
+Framingham 4,434 18 CVD (26.09%) Death (17.75%) 56.16%
+SUPPORT2 9,105 42 Cancer_Death (18.2%) Death_Other (49.9%) 31.9%
+Synthetic 30,000 12 * (25.33%) * (24.67%) 50.00%
 4.2 Experimental Setup
-The CRISP-NAM model is implemented using PyTorch.
-The experimental procedure is comprised of hyperparameter optimization followed by cross-validation evaluation phases.
-First, Optuna (Akiba et al., 2019) is employed to systematically search for optimal model configurations, tuning learning rate, $L_{2}$ regularization strength, dropout rates, network architecture (1-3 hidden layers with 32-256 units), and batch normalization settings using validation loss as the objective.
-The best configuration was then evaluated through stratified 5-fold cross-validation on multiple datasets previously described.
-Continuous features were normalized using standard scaling ($\mu=0,\sigma=1$).
-Missing values for categorical variables were imputed using mode imputation, and for continuous variables, mean imputation is applied for all datasets, except for the SUPPORT2 dataset for which median imputation was applied.
-Training employed the AdamW (Loshchilov and Hutter, 2017) optimizer to minimize the negative log-likelihood loss with a batch size of 256 and early stopping (patience=10) to prevent over-fitting.
+The CRISP-NAM model is implemented using PyTorch and the code is available from https://github.com/VectorInstitute/crisp-nam.
+We employed nested cross-validation to prevent data leakage during hyperparameter optimization and model evaluation.
+The approach consists of an outer 5-fold stratified cross-validation for performance evaluation and an inner 5-fold cross-validation for hyperparameter tuning within each outer fold.
+For each outer fold, Optuna (Akiba et al., 2019) is employed on the training partition to systematically search for optimal model configurations using the inner 5-fold cross-validation, tuning learning rate, $L_{2}$ regularization strength, dropout rates, network architecture (1-3 hidden layers with 8-128 units), and batch normalization settings using validation loss as the objective.
+The best configuration identified for each outer fold is then trained on the complete training partition and evaluated on the corresponding held-out test partition.
+Continuous features were normalized using standard scaling ($\mu=0,\sigma=1$) and categorical features were one-hot encoded.
+Missing categorical values were imputed using mode imputation.
+For continuous variables, mean imputation was used across all datasets.
+Training employed the AdamW (Loshchilov and Hutter, 2019) optimizer to minimize the negative log-likelihood loss with a batch size of 256 and early stopping (patience=10) to prevent over-fitting.
 Model performance was assessed using complementary metrics (Park et al., 2021) for discrimination and calibration.
@@
 $q_{.25}$ $q_{.50}$ $q_{.75}$ $q_{.25}$ $q_{.50}$ $q_{.75}$ $q_{.25}$ $q_{.50}$ $q_{.75}$
-FHS CRISP-NAM 1 0.850±.013 0.834±.009 0.812±.015 0.714±.022 0.712±.019 0.708±.019 0.059±.006 0.111±.004 0.153±.007
-2 0.785±.057 0.776±.029 0.769±.037 0.720±.033 0.719±.021 0.714±.021 0.040±.004 0.081±.006 0.130±.007
+FHS CRISP-NAM 1 0.843±.021 0.832±.028 0.811±.040 0.708±.022 0.700±.021 0.691±.023 0.249±.295 0.320±.273 0.353±.247
+2 0.793±.050 0.779±.027 0.771±.028 0.714±.029 0.713±.017 0.707±.014 0.041±.004 0.079±.006 0.127±.010
 NFG 1 0.678±.160 0.673±.149 0.666±.133 0.632±.150 0.629±.141 0.628±.134 0.058±.005 0.102±.005 0.134±.007
@@
 2 0.796±.053 0.779±.028 0.776±.031 0.737±.030 0.728±.018 0.724±.017 0.038±.003 0.072±.004 0.109±.005
-SUP CRISP-NAM 1 0.894±.024 0.839±.041 0.824±.052 0.682±.024 0.644±.017 0.636±.015 0.117±.075 0.216±.106 0.238±.111
-2 0.950±.037 0.879±.084 0.834±.107 0.869±.005 0.772±.005 0.734±.006 0.100±.007 0.170±.011 0.215±.018
+SUP CRISP-NAM 1 0.855±.065 0.802±.092 0.798±.100 0.665±.036 0.624±.029 0.617±.028 0.277±.268 0.380±.242 0.393±.227
+2 0.872±.188 0.838±.179 0.813±.175 0.804±.110 0.713±.097 0.680±.087 0.287±.266 0.308±.188 0.308±.129
 NFG 1 0.790±.036 0.777±.023 0.797±.020 0.620±.038 0.513±.025 0.438±.025 0.044±.003 0.100±.005 0.115±.005
@@
 2 0.955±.025 0.893±.069 0.860±.083 0.863±.008 0.745±.005 0.697±.006 0.087±.005 0.144±.006 0.171±.003
-PBC CRISP-NAM 1 0.986±.013 0.960±.029 0.953±.039 0.799±.016 0.783±.013 0.772±.009 0.126±.015 0.155±.016 0.176±.016
-2 0.977±.028 0.982±.015 0.985±.013 0.675±.020 0.662±.026 0.637±.019 0.314±.146 0.373±.152 0.388±.153
+PBC CRISP-NAM 1 0.988±.012 0.958±.028 0.942±.034 0.804±.019 0.785±.011 0.773±.006 0.131±.023 0.184±.042 0.242±.074
+2 0.952±.043 0.966±.023 0.972±.017 0.647±.009 0.635±.015 0.613±.016 0.469±.223 0.514±.217 0.537±.217
 NFG 1 0.853±.023 0.835±.011 0.824±.027 0.782±.017 0.765±.013 0.756±.016 0.143±.020 0.152±.008 0.170±.013
@@
 2 0.978±.036 0.983±.027 0.987±.020 0.728±.024 0.705±.019 0.690±.016 0.037±.006 0.056±.003 0.065±.004
-SYN CRISP-NAM 1 0.682±.011 0.675±.007 0.671±.011 0.585±.009 0.591±.005 0.587±.003 0.088±.014 0.188±.034 0.287±.047
-2 0.694±.021 0.677±.011 0.670±.009 0.594±.007 0.592±.008 0.586±.007 0.073±.016 0.160±.041 0.246±.059
+SYN CRISP-NAM 1 0.655±.019 0.660±.041 0.670±.081 0.551±.011 0.559±.008 0.555±.004 0.245±.336 0.301±.287 0.347±.219
+2 0.676±.026 0.653±.022 0.643±.021 0.569±.014 0.564±.014 0.560±.014 0.226±.246 0.356±.245 0.433±.172
 NFG 1 0.802±.011 0.771±.016 0.717±.016 0.741±.007 0.715±.008 0.670±.006 0.051±.001 0.096±.002 0.161±.003
@@
 While DeepHit generally achieves the highest performance, CRISP-NAM demonstrates competitive discrimination with the added benefit of interpretability through its additive structure.
-CRISP-NAM achieves discrimination metrics within 2-5% of DeepHit on clinical datasets, demonstrating its competitiveness.
-On FHS, the TD-AUC gap at $q_{0.25}$ is merely 0.004 (0.850 vs 0.854), with CRISP-NAM actually outperforming DeepHit at $q_{0.50}$ for Risk 1 (0.834 vs 0.831).
-Similarly, on SUPPORT2 Risk 1, CRISP-NAM exceeds DeepHit’s TD-AUC by substantial margins at $q_{0.25}$ (0.894 vs 0.779) and $q_{0.50}$ (0.839 vs 0.640).
-For PBC, while DeepHit achieves marginally higher TD-AUC, both models reach near-ceiling performance ($\geq 0.96$), making the practical difference negligible.
-The synthetic dataset represents the primary challenge, where DeepHit’s flexible architecture captures complex non-linear patterns that CRISP-NAM’s additive structure cannot fully represent.
-While DeepHit generally achieves lower Brier scores, CRISP-NAM maintains clinically acceptable calibration on most datasets.
-The calibration gap is smallest on FHS (within 0.01-0.02) and moderate on PBC Risk 1.
-The larger gaps on SUPPORT2 and PBC Risk 2 reflect the trade-off between interpretability and flexibility in probability estimation.
-CRISP-NAM exhibits lower variance across folds in several scenarios (e.g., FHS Risk 1: ±0.009 vs ±0.012 for DeepHit at $q_{0.50}$), suggesting more stable predictions.
-The performance gaps between CRISP-NAM and DeepHit are often clinically insignificant.
-A TD-CI difference of 0.02-0.03 rarely changes risk stratification decisions in practice.
-CRISP-NAM’s competitive performance, combined with its interpretability, positions it as a valuable alternative when model transparency is critical, particularly in high-stakes medical applications requiring explainable predictions.
+[…]
+Notably, CRISP-NAM’s calibration performance (Brier score) is also lower than NFG and DeepHit.
+This could be attributed to the loss function we implemented that optimizes for discrimination (ranking) rather than the calibration (probability accuracy).
+In addition, the event weighting could exacerbate the model’s lagging calibration.
+We will investigate improvements to the loss function to handle calibration more effectively in our future work.
 5.1 Interpretability Analysis
-We present shape function plots for the top 10 features per risk as learned by the CRISP-NAM model across 3 datasets: SUPPORT2, Framingham and PBC.
+Here, the shape function plots for the top 10 features per risk as learned by the CRISP-NAM model across 3 datasets: SUPPORT2, Framingham and PBC are presented.
 Each curve $s_{i,k}(x_{i})$ represents the marginal contribution of feature $x_{i}$ to the log cause-specific hazard for risk $k$.
@@
 Interpretation Note.
-These shape functions are associational, not causal, and may obscure interactions or time-varying effects.
+It should be noted that these shape functions are associational, not causal, and may obscure interactions between features.
 They are estimated under smoothness constraints and can extrapolate in regions with low data density, leading to amplified or flattened effects.
-As such, the shapes should be used primarily for model validation and debugging.
 Apparent patterns should be interpreted cautiously, corroborated on external cohorts, and discussed with domain experts before drawing scientific or clinical conclusions.
@@
 Figure 3: Shape functions computed with CRISP-NAM model for the top-10 important features: Framingham dataset
-Figure 2 presents shape functions from the CRISP-NAM model trained on the Framingham dataset, with separate plots for two competing risks: cardiovascular disease (Risk 1) and non–cardiovascular death (Risk 2).
-All functions are displayed on the log-hazard scale, where a unit increase of $+0.69$ corresponds to a doubling of the cause-specific hazard.
-For Risk 1, AGE, SYSBP, BMI, GLUCOSE, and the binary indicator DIABETES exhibit steadily increasing trends across most of their observed ranges.
-DIABP and AGE display similar steep upward slopes beyond their respective median values.
-TOTCHOL shows a mild U-shaped profile with a slight increase at high values.
-educ has an approximately linear decreasing trend across its range.
-The binary history flags PREVCHD and PREVAP show negative contributions to the log-hazard.Three data characteristics, rather than a real reversal of risk, explain this result:
+Figure 3 illustrates shape functions from the CRISP-NAM model trained on the Framingham dataset, with separate plots for two competing risks: cardiovascular disease (Risk 1) and non–cardiovascular death (Risk 2).
+All functions are displayed on the log-hazard scale, where a unit increase of +0.69+0.69 corresponds to a doubling of the cause-specific hazard.
+For Risk 1, GLUCOSE and BMI shows steep increase showing that higher blood glucose levels and increasing BMI are associated to higher risk of cardiovascular disease for this dataset.
+Shape plots reveal that the patients with diabetes, patients who are on medication for hypertension and who are current smokers (DIABETES, BPMEDS and CURSMOKE) have higher risk of cardiovascular disease which correlates well with known risk factors.
+Shape plot also reveals that women have lower risk of cardiovascular disease, compared to men and having more education is correlated with lower risk of cardiovascular disease.
+The binary history flags PREVCHD and PREVAP show negative contributions to the log-hazard.
+Three data characteristics, rather than a real reversal of risk, explain this result:
   1. 1.
@@
 Several observed patterns in the shape functions are consistent with known risk factors for cardiovascular and all-cause mortality.
-Established studies have linked increasing age, elevated systolic blood pressure, high glucose levels, and diabetes with heightened cardiovascular risk (Kannel and McGee, 1979; D’Agostino and Kannel, 1989).
+Established studies have linked elevated systolic blood pressure and high glucose levels with heightened cardiovascular risk (Kannel and McGee, 1979; D’Agostino and Kannel, 1989).
+Cigarettes per day displayed an exponential relationship with diminishing marginal effects at higher consumption levels, reflecting saturation of smoking-related harm pathways.
 The inverse trend for educational attainment aligns with literature on socioeconomic disparities in cardiovascular outcomes (Rosengren et al., 2019).
 The U-shaped relationship observed for BMI (Risk 2) has been previously noted in older adults and is often described as the “obesity paradox” (Amundson et al., 2010) as well as sparse data in the extreme ranges.
-Similarly, the positive association between elevated diastolic pressure and non-cardiovascular mortality may reflect constraints due to cohort truncation and sparsity in the extreme blood-pressure ranges.
+Additionally, previous cardiovascular conditions (angina pectoris and coronary heart disease) showed positive linear associations with mortality risk.
 Figure 4 shows how features contribute positively or negatively to the prediction.
@@
 Additionally, Figure 6 shows top 5 positive and top 5 negative important features that contributed to the model’s prediction.
-For Risk 1, age exhibits a monotonic increasing contribution to the predicted hazard.
-SGOT and alkaline display sharp increases followed by plateaus at higher values.
-The function for serBilir rises steeply at low values and saturates thereafter.
-serChol peaks at moderate levels and decreases subsequently.
-Binary indicators such as hepatomegaly_Yes and edema_edema despite diuretics show positive contributions.
-albumin demonstrates an inverted S-shaped curve: initially increasing, crossing zero near the median, and declining at higher values.
-platelet decreases steadily, with a minor upturn in the extreme right tail where data are sparse.
-For Risk 2, the shape for age decreases monotonically.
-SGOT rises rapidly and then flattens.
-histologic shows an approximately linear increasing trend.
-The shape for serBilir forms an inverted-U.
-albumin rises beyond the cohort median and levels off at higher values.
-ascites_Yes and hepatomegaly_Yes are associated with negative contributions.
-alkaline increases sharply before flattening.
-spiders_Yes exhibits a slight negative effect.
-Finally, serChol increases with higher values, especially in the upper tail, although data in that region are sparse.
-The rug plots accompanying each shape function reflect the distribution of the feature values and indicate regions with limited data support.These empirical patterns align with several well-established clinical insights in the context of Primary Biliary Cholangitis (PBC).
-For instance, older age, elevated liver enzymes such as SGOT and alkaline phosphatase, and increased serum bilirubin are recognized markers of disease severity and poorer prognosis (Dickson et al., 1989; Murillo Perez et al., 2023).
+For Risk 1, age exhibits a Sigmoid-shaped curve.
+alkaline displays sharp increases followed by plateaus at higher values.Biomarkers serBilir, serChol and prothrombin all show similar rises steeply at low values and plateus thereafter.
+Binary indicators such as hepatomegaly_Yes shows positive contributions.
+albumin demonstrates an inverted Sigmoid-shaped curve: initially increasing, crossing zero near the median, and declining at higher values.
+For Risk 2, the shape for age exhibits inverse sigmoidal shape with a decreasing trend as age increases signifying that transplantation risk decreases as patients progress with age.
+The shape for serBilir, alkaline and serChol rises steeply, before plateauing.
+These three elevated biomarkers indicate disease severity in PBC, which simultaneously increases both death risk and transplant priority.
+Examining the contribution scales for both risks for ascites_Yes shows a moderate negative contribution to Risk 1 ($\sim{-0.4}$) but a much stronger negative contribution to Risk 2 ($\sim{-1.0}$).
+From a model validation perspective, this suggests the model has learned that ascites presence is associated with reduced likelihood of both outcomes, but particularly transplantation.
+The asymmetric magnitudes indicate the model distinguishes between the two competing risks rather than simply treating ascites as a general severity marker.
+The rug plots accompanying each shape function reflect the distribution of the feature values and indicate regions with limited data support.
+These empirical patterns align with several well-established clinical insights in the context of Primary Biliary Cholangitis (PBC).
+For instance, older age, elevated liver enzymes such as alkaline phosphatase, and increased serum bilirubin are recognized markers of disease severity and poorer prognosis (Dickson et al., 1989; Murillo Perez et al., 2023).
 The decreasing transplant hazard with age may reflect clinical prioritization criteria that favour younger candidates for organ allocation.
 The non-linear shape for albumin aligns with its known role as a proxy for liver synthetic function, where low levels indicate hepatic decompensation.
-Histologic stage progression, from fibrosis to cirrhosis, is a standard determinant in transplant eligibility, consistent with the monotonic rise observed in its shape function.
+Histologic stage (histologic) progression, from fibrosis to cirrhosis, is a standard determinant in transplant eligibility, consistent with the monotonic rise observed in its shape function.
 Additionally, ascites and hepatomegaly are classical signs of advanced liver disease, often associated with higher mortality and reduced transplant suitability.
-The distinct inverted-U pattern for bilirubin in the transplant hazard may reflect clinical decision-making thresholds, while the rising contribution of cholesterol could be interpreted in light of its association with bile duct dysfunction in PBC.
+Edema in PBC patients indicates advanced liver disease.
+In the competing risks framework, patients with edema have higher disease severity and thus receive higher priority for transplantation due to their urgent medical need.
+The positive contribution to transplantation risk reflects the clinical reality that transplant allocation systems prioritize sicker patients and those with edema are more likely to receive transplants because it serves as a marker of advanced disease requiring urgent intervention.
 5.1.3 SUPPORT2 Dataset
@@
 Figure 7 presents feature importance and shape functions from the CRISP-NAM model trained on the SUPPORT2 dataset, which distinguishes cancer-specific mortality (Risk 1) from death due to other causes (Risk 2).
-All shape functions are plotted on the log-hazard scale.
-For Risk 1, the variable sps shows a near-linear increase, indicating a progressive rise in the predicted hazard across its range.
-The shape function for age follows an inverted-U profile, with risk increasing up to approximately 65 years and declining thereafter.
-Binary indicators dementia_1 and dnr_no_dnr exhibit negative contributions to the log-hazard.
-The shape for avtisst is flat at lower durations of mechanical ventilation but shows a steep increase beyond approximately 3 days.
-Negative contributions for dnr_no_dnr and dementia_1 may reflect treatment-limiting decisions or documentation timing effects, as suggested in earlier SUPPORT studies (Hakim et al., 1996).
-For Risk 2, the severity scores aps and sps both show sharp increases followed by plateaus.
+We highlight several observations from the shape plots for both risks.
+For Risk 1, the shape function for age shows a lower risk of death from cancer for younger patients, then steadily increasing risk of cancer-related death up to approximately 65 years and declining risk for very old patients.
+The binary indicator race_black exhibits a slight negative contribution to the log-hazard for Risk 1 showing that black patients appear to have lower mortality due to cancer.
+The apparent protective effect of Black race contradicts established epidemiological evidence (Saka et al., 2025) demonstrating higher cancer mortality rates in Black population.
+This could indicate insufficient sample representation or selection biases inherent to the SUPPORT2 dataset.
+Similarly, the inverse relationship between number of comorbidities and cancer death risk suggests competing mortality mechanisms, where patients with multiple comorbidities may succumb to other medical conditions before cancer progression becomes the primary threat.
+The shape for avtisst is lower durations of mechanical ventilation but shows a steep increase beyond certain number of days.
+pafi (oxygen ratio) exhibits a spike at very low values ($\sim{0.7}$), then drops to steady negative contribution ($\sim{0.3}$).
+Other biomarkers such as sodium sod shows an inverted U-shape peaking near normal levels, suggesting both low sodium levels as well as very high sodium levels increases risk.
+For Risk 2, the severity scores aps show sharp increases followed by plateaus.
 The avtisst variable again displays a marked rise in hazard for durations exceeding 3 days.
 Cost-related variables charges and totcst show decreasing trends.
-The binary indicator dnr_no_dnr also shows a downward slope across its two levels.
-Rising sps and aps scores are consistent with the role of physiological instability and organ dysfunction in predicting mortality (Knaus et al., 1995).
+The binary indicator dnr_no_dnr shows that patients without DNR (Do Not Resuscitate) orders (dnr_no_dnr = 1) demonstrated substantially lower contributions to non-cancer death risk compared to those with DNR orders present.
+This pattern aligns with clinical expectations, as DNR orders typically indicate patients with poor overall prognosis, advanced chronic diseases, or end-stage conditions who are at elevated risk for cardiovascular, respiratory, or multi-organ failure.
+Rising aps scores are consistent with the role of physiological instability and organ dysfunction in predicting mortality (Knaus et al., 1995).
 The inverted-U for age in cancer mortality likely reflects competing risks from other causes in older individuals (Hayes-Larson et al., 2020).
-The association between prolonged ventilation (avtisst $>3$ days) and increased hazard aligns with the known severity of illness in patients requiring extended respiratory support.
+The association between prolonged ventilation (avtisst >3>3 days) and increased hazard aligns with the known severity of illness in patients requiring extended respiratory support.
 Cost variables likely act as proxies for length of stay or illness trajectory rather than direct predictors.
@@
 CRISP-NAM is particularly valuable in high-stakes healthcare ML applications requiring mechanistic understanding such as investigating associational relationships, assessing treatment efficacy, or designing targeted interventions, especially when competing events represent distinct clinical processes (Austin et al., 2016).
+We acknowledge that CRISP-NAM inherits the proportional hazards assumption from the Cox framework, requiring that covariate effects on each cause-specific hazard remain constant over time.
+This assumption may be violated when covariate effects vary temporally, such as biomarkers having different predictive power for early versus late events.
+There are two possible avenues for future work: (i) exploration of temporal FeatureNets capable of learning how feature contributions evolve over time.
+The caveat that is that this approach would likely increase computational complexity during training and potentially demand larger datasets.
+Additionally, we will investigate using a modified loss function that takes calibration (Brier Score) into account.